OpenStudy (anonymous):

y = 1/2cotπx. state the period

OpenStudy (anonymous):

well like as u know the periond is 360/k or 2pi/k

OpenStudy (anonymous):

well here our k is pi so we divide 2pi/pi=2 so our period is 2. liek i just guessed this and it kinda makes sense

OpenStudy (anonymous):

in terms of pi i guess

OpenStudy (anonymous):

what do u mean in terms of pi?

OpenStudy (anonymous):

like our period wld be 2 radians

OpenStudy (anonymous):

if u want degrees like we can do that tooo

OpenStudy (anonymous):

that wld 360/pi=115 degrees approx

OpenStudy (anonymous):

how about this one? y = tanx/2

OpenStudy (anonymous):

plz put barackets so i can like it makes it clearer

OpenStudy (anonymous):

y = tan(x/2)

OpenStudy (anonymous):

ooook that is better ok but btw i just dbl checked and 115degrees is teh same thing as 2 radians obv so liek whtvr anser u give just specify if it is radians or degrees

OpenStudy (anonymous):

btw do u have to give ur answer is radians or degrees

OpenStudy (anonymous):

does not say.

OpenStudy (anonymous):

ok so here our k is =(1/2) so we do 360/(1/2) which is essence 360*2 do u see that?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):